OpenStudy (anonymous):

Consider an x-ray tube with a Cu anode and a monochromator consisting of a nickel single crystal with (111) planes parallel to the surface of the crystal. What is the lowest voltage capable of producing x-rays through a monochromator?

OpenStudy (anonymous):

3053

OpenStudy (aaronq):

lol how do you even do this?

OpenStudy (anonymous):

λ = 2diii sinθ => λmax=2diii θ=90 eV = hc/λsw => smallest voltage => longest wavelength capable of diffraction \[diii=\frac{ a }{ \sqrt{1^{2}+1^{2}+1^{2}}} \] α = 3.53*10^-8 cm \[d_{iii}=2.04*10^{-8} cm\] \[λ_{\max}=4.07*10^{-8}cm\] \[V=\frac{ \left( 6.626*10^{-34} \right)\left( 3*10^{8} \right) }{ \left( 1.6*10^{-19} \right)\left( 4.07*10^{-10} \right) } = 3053 V\]